If g(y) = y/(1-y), show that 1/2[g(y) + g(-y)]= g(y^2).

1
Expert's answer
2022-06-12T14:57:03-0400
"g(y)=\dfrac{y}{1-y}, y\not=1"

"g(-y)=\dfrac{-y}{1-(-y)}, y\not=-1"

"\dfrac{1}{2}[g(y)+g(-y)]=\dfrac{1}{2}[\dfrac{y}{1-y}+\dfrac{-y}{1+y}]"

"=\dfrac{1}{2}(\dfrac{y+y^2-y+y^2}{1-y^2})=\dfrac{y^2}{1-y^2}"

"=g(y^2), y\not=\pm1"

Therefore


"\dfrac{1}{2}[g(y)+g(-y)]=g(y^2), y\not=\pm1"